OpenStudy (anonymous):

limit as x approaches infinity of 1-cosx/x^2 using sandwich theorem

OpenStudy (anonymous):

i just have to say sandwich is a funny name.

OpenStudy (anonymous):

that was helpful..

OpenStudy (anonymous):

sorry...

OpenStudy (anonymous):

since \(-1\leq\cos(x)\leq 1\) we know \(-2\leq 1-\cos(x)\leq 0\)

OpenStudy (anonymous):

ok my inequality stinks, let me try again

OpenStudy (anonymous):

\[0\leq 1-\cos(x)\leq 2\] that looks better

OpenStudy (anonymous):

where'd you get the second inequality?

OpenStudy (anonymous):

which means can "sandwich" \[0\leq \frac{1-\cos(x)}{x^2}\leq \frac{2}{x^2}\]

OpenStudy (anonymous):

mmmm sandwich oh first inequality is wrong

OpenStudy (anonymous):

\[-1\leq \cos(x)\leq 1\] is right and so \(1-\cos(x)\) is smallest when \(\cos(x)=1\) in which case you get 0, and largest when \(\cos(x)=-1\) in which case you get 2

OpenStudy (anonymous):

pretty clear that as \(x\to \infty\) we know \(\frac{2}{x^2}\to 0\) so you have a zero on the left and a zero on the right and that is your sandwich

OpenStudy (anonymous):

wait I don't understand how cos(x) = 1 you get 0? and such?

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

the largest cosine can be is 1, so the smallest -cosine can be is -1 and so the smallest 1-cosine can be is 0

OpenStudy (anonymous):

this is really messing me up, if this is only the beginning of calculus, i'm scared

OpenStudy (anonymous):

i think you are fretting for very little reason. calc can be difficult, but this part is just reasoning, not computation is it clear that the smallest 1 - cosine can be is 0 ?

OpenStudy (anonymous):

it is because cosine is stuck between minus one and one, and therefore so is - cosine

OpenStudy (anonymous):

it's not making sense.. 1 - (-1) is 2....

OpenStudy (anonymous):

yes exactly, that is the very largest \(1-\cos(x)\) can be the biggest it can be is if cosine is minus one, and in that case you get 2 that is the maximum it can be

OpenStudy (anonymous):

in other words \(1-\cos(x)\) cannot bet greater than 2 you have just shown it

OpenStudy (anonymous):

so cos(x) can't be greater than 1, but that doesn't mean 1-cos(x) can't be greater than 1

OpenStudy (anonymous):

yes you are right. it doesn't mean \(1-\cos(x)\) cannot be greater than one, it means \(1-\cos(x)\) cannot be greater than 2 as you said above

OpenStudy (anonymous):

okay so \[1-\cos(x) \le 2\]

OpenStudy (anonymous):

yes for sure. it is more or less clear right?

OpenStudy (anonymous):

alright, continue, it's catching on

OpenStudy (anonymous):

perhaps the 1 at the beginning is confusing suppose instead we had \(10-\cos(x)\) what is the biggest that could be? answer is 11, because of the bound on the cosine what is the smallest? answer is 9 for the same reason

OpenStudy (ujjwal):

I didn't know there was a sandwich theorem! Thanks @satellite73 and @jamroz

OpenStudy (anonymous):

in any case we see that the very largest \(1-\cos(x)\) can be is 2 what is the smallest it could be? answer is 0. it would be zero if \(\cos(x)=1\) you would get \(1-1=0\) and it cannot be any less than 0, because cosine cannot be any larger than 1

OpenStudy (anonymous):

okay that makes sense, understanding that, what do I do now?

OpenStudy (anonymous):

not that hold on, new paragraph to read

OpenStudy (anonymous):

ok let me know when we continue

OpenStudy (anonymous):

okay got it, now what?

OpenStudy (anonymous):

thank you so much for your patience by the way, really appreciated

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

ok now we almost have our sandwich. \(1-\cos(x)\) is stuck between 0 and 2, so we write the inequality \[0\leq 1-\cos(x)\leq 2\]

OpenStudy (anonymous):

i have that down, now what?

OpenStudy (anonymous):

then dividing by \(x^2\) gives \[0\leq \frac{1-\cos(x)}{x^2}\leq \frac{2}{x^2}\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

understood, proceed

OpenStudy (anonymous):

then we take the limit as \(x\to \infty\)

OpenStudy (anonymous):

on the left we get 0 because there isn't even an \(x\) in it, it is just always 0

OpenStudy (anonymous):

and what about \[\lim_{x\to \infty}\frac{2}{x^2}\]? well that has to be 0 as well, because \(x^2\) is getting larger and larger, but the numerator is stuck at 2

OpenStudy (anonymous):

I don't understand why it has to be 0?

OpenStudy (anonymous):

oops i meant \[\lim_{x\to \infty}\frac{2}{x^2}=0\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} 2\div x^{2}\]

OpenStudy (anonymous):

why does that have to be 0?

OpenStudy (anonymous):

couldn't it still be 1 and work?

OpenStudy (anonymous):

x is getting larger and larger. think of a number with a huge denominator and a 2 in the numerator that number is going to be very close to 0 for example if \(x=10\) then \(x^2=100\) and \(\frac{2}{100}=.02\) and that is only if \(x=10\) not a very large number now suppose \(x=1000\) then \(x^2=1,000,000\) and \(\frac{2}{x^2}=0.000002\) getting very close to 0

OpenStudy (anonymous):

but if x = 1 then x^2 = 1 and it'd be 2/1 = 2?

OpenStudy (anonymous):

brb, away from the keyboard for a minute

OpenStudy (anonymous):

you can make it as close to zero as you like, by making x larger and larger the bigger x gets, the smaller \(\frac{2}{x^2}\) gets

OpenStudy (anonymous):

oh yes, at 1 you get two but you are taking the limit as x goes to infinity, not 1 x is getting bigger and bigger, not closer and closer to 1

OpenStudy (anonymous):

true, true. you make your point. so that's how I validate it's going to 0 using sandwich?

OpenStudy (anonymous):

i'm trying one for the lim x as it approaches -infinity for sinx/x and i'm using the sandwich kind of.. is this the right path? \[-1 \le sinx \le 1\] \[\frac{ -1 }{ x } \le \frac{ sinx }{ x } \le \frac{ 1 }{ x }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then where would I go from there? it's catching on thank you so so much

OpenStudy (anonymous):

to answer your question above, since the limit on the left is 0 and the limit on the right is 0, the limit in the middle must be 0 as well

OpenStudy (anonymous):

\[0\leq \lim_{x\to \infty}\frac{1-\cos(x)}{x^2}\leq 0\] indicates that the middle one must be zero as well

OpenStudy (anonymous):

that makes complete sense! :D

OpenStudy (anonymous):

\[\frac{ -1 }{ x } \le \frac{ \sin(x) }{ x } \le \frac{ 1 }{ x }\] just like the previous one take the limit as \(x\to \infty\) on the right and the left you will get 0 in both places

OpenStudy (anonymous):

you will have to take my word for it that soon this will all seem completely obvious since sine is stuck between -1 and 1, as the \(x\) in the denominator gets larger and larger the whole thing must go to zero

OpenStudy (anonymous):

because with -1/x when x gets larger it gets closer and closer to 0 as it approaches -infinity right?

OpenStudy (anonymous):

yes for sure

OpenStudy (anonymous):

the minus one up top is a red herring. it could be 57,208 and it would still go to zero as the denominator goes to minus infinity

OpenStudy (anonymous):

but isn't sinx/x = 1?

OpenStudy (anonymous):

that is the limit as x goes to zero, not as x goes to infinity

OpenStudy (anonymous):

\[\lim_{x\to 0}\frac{\sin(x)}{x}=1\] but \[\lim_{x\to \infty}\frac{\sin(x)}{x}=0\]